Nancy jogs and walks to school each day. She averages 4 Km/h walking and 8 Km/h jogging.The distance from home to school is 6 Km, and she makes the trip in 1hour.How far does she jog in a trip?
Answer by mananth(16946)

Nancy jogs and walks to school each day. She averages 4 Km/h walking and 8 Km/h jogging.The distance from home to school is 6 Km, and she makes the trip in 1hour.How far does she jog in a trip?
..
walking 4km/h
jogging 8km/h
..
distance = 6 km to school
time taken = 1hour
..
jogging distance be x km
walking distance = 6-x
..
t = d/r
time jogging + time walking = 1
x/8 + x/4 = 1
(x+2x)/8 = 1
multiply by 8
3x = 8
/3
x= 8/3
x=2.66 km
...
m.ananth@hotmail.ca